OpenStudy (anonymous):

having trouble isolating t... 3t^2 + 12t = s

OpenStudy (anonymous):

Hello! Sorry you could not get any help! :( Here is a video for your future questions like this. https://www.khanacademy.org/math/algebra/solving-linear-equations-and-inequalities/solving_for_variable/v/rearrange-formulas-to-isolate-specific-variables

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ 3t^2+12t=s \hspace{.33em}\\~\\ t^2+4t=\dfrac{s}{3} \hspace{.33em}\\~\\ t^2+2(2t)+2^2=\dfrac{s}{3}+2^2 \hspace{.33em}\\~\\ (t+2)^2=\dfrac{s+12}{3} \hspace{.33em}\\~\\ t=\pm\sqrt{\dfrac{s+12}{3}}-2 \hspace{.33em}\\~\\ }\end{align}\)